what is the mean for the combined set if the first sample has n=3 and the second sample has n=6
1
Expert's answer
2019-09-16T10:49:10-0400

The combined mean for two sets will depend on also the means for the each respective set:

"x_{a+b} = \frac{n_ax_a+n_bx_b}{n_a+n_b}" ,

where:

"x_a" = the mean of the first set,

"n_a" = the number of items in the first set,

"x_b" = the mean of the second set,

"n_b" = the number of items in the second set,

"x_{a+b}" = the combined mean.


Using the information we have,

"n_a = 3; n_b = 6, \\
x_{a+b} = \frac{1}{9}(3x_a+6x_b) = \frac{1}{3}(x_a+2x_b)."